Abel Company uses activity-based costing. The company has two products: A and B. The annual production and sales of Product A is 200 units and of Product B is 400 units. There are three activity cost pools, with estimated costs and expected activity as follows: 



Expected Activity

Activity Cost PoolEstimated CostProduct AProduct BTotal
Activity 1$16,660600100700
Activity 218,4501,1007001,800
Activity 39,73160160220

22. The cost per unit of Product A is closest to which of the following? 

A. $27.91. 

B. $56.38. 

C. $141.04. 

D. $70.52.
Answer #1

Activity rate=Respective cost/Respective activity

Activity rate for:

Activity 1=(16,660/700)=$23.8

Activity 2=(18450/1800)=$10.25

Activity 3=(9731/220)=$44.2318182(Approx)

Hence cost for A=(23.8*600)+(10.25*1100)+(44.2318182*60)

=$28208.91(Approx)

Hence cost per unit=28208.91/200

=$141.04(Approx)